OpenStudy (nightmarenight):

If f(x) varies directly with x^2, and f(x) = 75 when x = 5, find the value of f(4).

OpenStudy (jdoe0001):

\(\bf \begin{array}{cccllll} \textit{something }&\textit{varies directly to }&\textit{something else}\\ \quad \\ \textit{something }&={\color{red}{ \textit{some value }}}&\textit{something else}\\ \quad \\ y&={\color{red}{ n}}&x&\implies y={\color{red}{ n}}x \end{array}\\ \quad \\ f(x)=y={\color{red}{ n}}x^2\qquad f(x)=y = 75\qquad x = 5\implies 75={\color{red}{ n}}5^2\) solve for "n" once you get "n", to get f(4), recall that \(\bf f(4)=({\color{red}{ n}})4^2\)

OpenStudy (nightmarenight):

Is there more? I'm not quite following @jdoe0001

OpenStudy (jdoe0001):

hmmm what did you get for "n" anyway?

OpenStudy (nightmarenight):

That's what I'm uncertain about, how to solve for "n"

OpenStudy (jdoe0001):

\(\bf (x)=y={\color{red}{ n}}x^2\qquad f(x)=y = 75\qquad x = 5\implies 75={\color{red}{ n}}5^2\\ \quad \\ 75=5^2\cdot n\) so.... there.. notice it's just a simple linear equation

OpenStudy (jdoe0001):

what do you think is "n"?

OpenStudy (nightmarenight):

uh. I don't follow?????? @jdoe0001

OpenStudy (jdoe0001):

\(\bf 75=5^2\cdot n\implies 75=25\cdot n \) have you done any linear equations yet?

OpenStudy (nightmarenight):

no, this is my first time doing these kinds of equations, how do i solve for n? @jdoe0001

OpenStudy (jdoe0001):

if you divide both left-side and right-side by 25, what would it give you?